# One-Shot Federated Learning with Classifier-Free Diffusion Models

### v1

@@ Abstract
Federated learning (FL) enables collaborative learning without data centralization but introduces significant communication costs due to multiple communication rounds between clients and the server. One-shot federated learning (OSFL) addresses this by forming a global model with a single communication round, often relying on the server’s model distillation or auxiliary dataset generation - often through pre-trained diffusion models (DMs). Existing DM-assisted OSFL methods, however, typically employ classifier-guided DMs, which require training auxiliary classifier models at each client, introducing additional computation overhead. This work introduces OSCAR (One-Shot Federated Learning with Classifier-Free Diffusion Models), a novel OSFL approach that eliminates the need for auxiliary models. OSCAR uses foundation models to devise category-specific data representations at each client, seamlessly integrated into a classifier-free diffusion model pipeline for server-side data generation. OSCAR is a simple yet cost-effective OSFL approach that outperforms the state-of-the-art on four benchmarking datasets while reducing the communication load by at least $\mathbf{99}\%$ [^1] https://anonymous.4open.science/r/oscar-FE04/..
Index Terms:
Federated Learning, One-Shot Federated Learning, Diffusion Models
@@ I Introduction
Federated Learning (FL)  [1] is a decentralized machine learning (ML) training methodology that enables multiple clients to collaboratively train a global model without moving the data to a central location, addressing the concerns on data privacy and ownership in the age of growing data privacy regulations. This has led to the application of FL to various domains, including autonomous vehicles  [2], the Internet of Things (IoT)  [3], and healthcare  [4]. However, since the participating clients usually own nonindependent and identically distributed (non-IID) data, an optimal global model is learned through multiple communication rounds between the clients and the server, causing high communication overhead  [5]. Several strategies, including client selection  [6], update compression  [7], and update dropping  [8, 9], have been proposed to reduce communication load in each communication round. However, these strategies still require client synchronization and suffer under non-IID data distribution across clients, as local client models can drift from the global model at each communication round  [10].
@@
Figure 1: The number of uploaded parameters by each client and accuracy for various algorithms on OpenImage dataset and ResNet-18.
@@
One-shot federated learning (OSFL)  [11, 12] offers an alternative, where the global model is learned through a single communication round between the clients and the server. OSFL can also reduce the impact of heterogeneous data distribution as the global model is not directly formed from the local models. The single communication round can further benefit in scenarios where FL suffers from client dropout or stragglers (i.e., slow-communicating clients)  [13]. Existing OSFL approaches  [14, 15, 16] rely on auxiliary dataset generation or knowledge distillation to form the global model. Knowledge distillation approaches usually require an auxiliary public dataset as the knowledge transfer medium  [15]. On the other hand, dataset generation methods employ pre-trained generative models (e.g., diffusion models) to generate new data for training the global model. DMs are pre-trained with vast amounts of data, and with proper guidance, these DMs can generate realistic images that resemble a desired distribution. DMs can have an immense impact on FL, as new data resembling the clients’ distribution can be generated without access to the raw dataset of the participating clients.
@@
Current studies incorporating diffusion models in OSFL  [13, 17, 16, 18] utilize classifier-guided DMs. Employing classifier-guided DMs has two potential disadvantages: they require auxiliary classifier training at each client, introducing computational and communication burdens at each client. Furthermore, in some cases, the diffusion model needs to be downloaded to the clients [16]. Classifier-free DMs  [19] solve these challenges by integrating the conditioning directly into the model, which is also adopted by most of the prevalent image generative models [20, 21]. FMs  [22, 23] can be employed for the encoding generation without training or fine-tuning. Replacing the classifier models with FMs significantly reduces the client upload size compared to classifier-guided DM-based OSFL approaches, as shown in Fig 1. The seamless integration of pre-trained FMs and DMs in OSFL simplifies the overall framework, reduces communication load, and enhances scalability and efficiency across heterogeneous client datasets.
@@ II-A One-Shot Federated Learning
Existing OSFL approaches can be divided into two categories based on their methodology. The first category utilizes knowledge distillation to learn a global model through either data distillation [15] or model distillation [24]. In distilled one-shot federated learning (DOSFL)  [15], the clients share distilled synthetic data with the server, which is utilized for the global model training. FedKT  [24] utilizes a public auxiliary dataset and student-teacher models trained at clients to learn a global student model. The second category of methods uses auxiliary data generation at the server based on intermediary information shared by the clients. DENSE  [14] trains a generator model on local classifiers, later used to generate auxiliary data for global model training. In FedCVAE  [25], the server aggregates the decoder part of the conditional variational encoders (CVAE) trained at each client and generates auxiliary data for the global model. FedDiff  [17] aggregates locally trained diffusion models for forming a global diffusion model for data generation. FedCADO  [13] utilizes classifiers trained at each client to generate data for global model training via classifier-guided pre-trained diffusion models (DMs). FedDISC  [26] utilizes data features for data generation via pre-trained DMs.
@@ II-B Federated Learning with Foundation Models
The emergence of foundation models (FMs), both large language models (LLMs)  [23] and vision language models (VLMs)  [22], has impacted the landscape of machine learning. The application of these FMs, however, has been understudied in FL. Yu et al.,  [27] and Charles et al.,  [28] explore training FMs in FL setups. PromptFL  [29] investigates prompt learning for FMs under data scarcity in FL settings. FedDAT  [30] proposes a federated fine-tuning approach for multi-modal FMs. FedPCL  [31] integrates FMs into the traditional FL process to act as class-wise prototype extractors. While FMs have the potential to mitigate data heterogeneity and communication load in FL, their full potential has not been utilized in FL settings.
@@ III-A Diffusion Models
$\mathcal{L}_{t}(\theta)=\mathbb{E}_{z_{0}\sim q(\mathcal{E}(x_{0})),\;\epsilon\sim\mathcal{N}(0,\;\mathbf{I}),\;t}\left[\left\|\epsilon-\hat{\epsilon}_{t}\right\|^{2}\right],\;\hat{\epsilon}_{t}=\epsilon_{\theta}(x_{t},\;t,\;y),$ (3)
where $\hat{\epsilon}_{t}$ provides an estimate of the score function used for data generation during the reverse process.
@@ IV OSCAR: One Shot Federated Learning with Classifier-Free Diffusion Models
Traditional FL setup consists of a central server and a set of clients $\mathcal{R}$, where each client $r\in\mathcal{R}$ trains a local model $\mathbf{w}_{r}$ on its local dataset $\mathcal{D}_{r}$ in each iteration and communicates it to the server. The server is responsible for forming a global model $\mathbf{w}$ from the local client models through an aggregation function like federated averaging (FedAvg). Considering the intrinsic data heterogeneity in FL, traditional FL algorithms require multiple communication rounds between the clients and the server to form a global model, leading to significant communication overhead.
To reduce the communication load in the FL setup, we propose OSCAR, a novel one-shot federate learning (OSFL) approach. OSCAR integrates foundation and pre-trained classifier-free generative models, specifically Stable Diffusion [21], and learns a global model from clients’ category-specific representations. OSCAR facilitates global model learning within a single communication round under non-IID data distribution among the clients. As illustrated in Fig. 2, the OSCAR pipeline can be divided into four steps: (1) generating descriptions of the client’s client-specific data, (2) encoding features from the generated descriptions, and transmitting them to the server (3) transmitting the category-specific representations to the server, and (4) generating data on the server to facilitate final model training.
@@
Unlike existing DM-assisted OSFL approaches, OSCAR eliminates classifier training and utilizes the clients’ category-specific data features as conditioning for the diffusion model (DM). Each client follows a two-step approach to generate category-specific encodings. First, the client uses a vision-language model (VLM), specifically BLIP [35], to generate textual descriptions for all their images. Then, the client uses a text encoder, specifically CLIP [36] to generate category-specific text-encodings from the textual descriptions, as shown in Eq. 6. A classifier-free diffusion model can utilize the CLIP encodings directly as text conditioning to generate new data.
@@
$x_{T-1}=\frac{1}{\sqrt{\alpha_{T}}}\left(\sqrt{\alpha_{T}}x_{T}-\hat{\epsilon}_{T}(\bar{y}_{c})\right)+\sigma_{T}\mathcal{N}(0,\mathbf{I}),$ (9)
[…]
@@
After generating the global synthesized dataset $\mathcal{D}_{syn}$, the server trains a centralized model $\mathbf{w}$, specifically a ResNet-18 classifier, on the synthesized dataset. This approach not only reduces the dependency on client synchronization and availability at each training and communication round but also ensures that the model can generalize well to non-IID data across clients. The server communicates the global model $\mathbf{w}$ to all clients after formation, to be later used for inference locally.
@@ V Experimental Setup
  • NICO++  [37] The dataset contains images of size 224 $\times$ 224 from 60 different categories, where each category contains images from different domains. The dataset is designed to evaluate models under heterogeneous environment settings. The dataset has two settings. In Common NICO++, all the categories share the same six domains: [autumn, dim, grass, outdoor, rock, and water]. On the other hand, the Unique NICO++ contains different domains for each category.
  • DomainNet  [38] The DomainNet dataset contains images of 345 categories over six domains. This work uses a subset of the DomainNet dataset with 90 categories.
  • OpenImage  [39] OpenImage is a multi-task dataset with over 1.7 million images across 600 categories. This work uses a subset of 120 categories from the dataset following the pre-processing in  [13].
@@
The data has been divided among all parties in a nonindependent and identically (non-IID) manner. Specifically, the data is feature distribution skewed, as each client owns data about a single domain from each category in the NICO++ and DomainNet datasets. For Openimage, the classes have been divided into six similar subgroups, where each client owns a single category from each subgroup. The number of clients is fixed to six, aligning with the number of domains in all the datasets.
@@
We have considered local learning, federated learning, and state-of-the-art DM-assisted OSFL approaches for comparison against OSCAR. Each client trains a local standalone model on its data in local training. FedAvg [1], FedProx [40], and FedDyn [41] consider traditional FL setups with minor variations in the local objective and aggregation functions. FedDISC [26] and FedCADO [13] are DM-assisted OSFL approaches and train auxiliary models for image generation at the server.
@@ VI Results and Analysis
To facilitate direct comparison with existing baseline approaches, the classifier network in the main results reported is a ResNet-18. However, the synthesized data has more potential that advanced backbones can utilize. The results for NICO++ Unique and NICO++ Common datasets with different classifier networks are reported in Table II. These results indicate that the generated data can potentially improve the global model’s optimality with an improved model architecture and may even improve more as the number of images per category increases. In the ResNet family, ResNet-101 performs the best, while the base version of the vision transformer (ViT B-16) has the best overall performance. The number of images per category for each client in these experiments is set to 10.
TABLE II: Accuracy (in %) on the test set for OSCAR with different classifier networks at the server.
@@
Number of generated images
This section examines the impact of generated dataset size on the global model performance in OSCAR. Traditionally, the increase in dataset size impacts the performance positively. In this case, while in the initial increase in synthesized dataset size boosts the model performance, the performance remains constant, or in some cases decrease, after a certain threshold. This can also mean that diffusion model synthesized data can act as auxiliary for training OSFL approaches rather than a replacement. Table III shows the result for OSCAR with different number of samples synthesized per category of each client.
@@
Communication Analysis
As OSCAR only uploads the data encodings, it uploads the least parameters from each client. OSCAR uploads less than 1% of the number of parameters compared to the SOTA models. OSCAR achieves this by eliminating the classifier training, and hence classifier uploading, and each client only communicates 512 parameters for each category. Table IV shows the number of parameters each client uploads. FedCADO trains a classifier model; hence, each client uploads a model with 11.69 million parameters. While FedDISC reduces the communication size by more than 60% compared to FedCADO at each client, the upload size from each client is still 100 times higher than OSCAR. The difference becomes more prominent as the number of clients increases.
@@ VII Conclusion
In this work, we propose OSCAR, a novel one-shot federated learning approach that utilizes pre-trained vision language models and classifier-free diffusion models (DMs) to train a global model in a single communication round in FL settings. OSCAR eliminates the need for training a classifier model at each client by replacing the classifier-guided DM with a classifier-free DM in the image synthesis phase. OSCAR generates category-specific data representations for each client through BLIP and CLIP foundation models, which are communicated to the server. The server generates new data samples and trains a global model on the generated data. OSCAR reduces the communication load by reducing the client upload size by 100X, compared to state-of-the-art DM-assisted OSFL approaches, while exhibiting superior performance on four benchmarking datasets.
In future work, we want to extend OSCAR to fuse the knowledge of auxiliary generated datasets with existing learned knowledge at each client. Existing DM-assisted OSFL approaches neglect the local models after data synthesis at the server, as it leads to computational overhead for the clients. A fine balance between both can lead to a more efficient global model.

### v2

@@ Abstract
Federated learning (FL) enables collaborative learning without data centralization but introduces significant communication costs due to multiple communication rounds between clients and the server. One-shot federated learning (OSFL) addresses this by forming a global model with a single communication round, often relying on the server’s model distillation or auxiliary dataset generation - mostly through pre-trained diffusion models (DMs). Existing DM-assisted OSFL methods, however, typically employ classifier-guided DMs, which require training auxiliary classifier models at each client, introducing additional computation overhead. This work introduces OSCAR (One-Shot Federated Learning with Classifier-Free Diffusion Models), a novel OSFL approach that eliminates the need for auxiliary models. OSCAR uses foundation models to devise category-specific data representations at each client which are integrated into a classifier-free diffusion model pipeline for server-side data generation. In our experiments, OSCAR outperforms the state-of-the-art on four benchmark datasets while reducing the communication load by at least $\mathbf{99}\%$ [^1] https://github.com/obaidullahzaland/oscar.
@@ I Introduction
Federated Learning (FL)  [20] is a decentralized machine learning (ML) training methodology that enables multiple clients to collaboratively train a global model without moving the data to a central location, thus addressing concerns about data privacy and ownership in the age of growing data privacy regulations. This has led to the application of FL to various domains, including autonomous vehicles  [31], the Internet of Things (IoT)  [30], and healthcare  [3]. However, since the participating clients usually own data that is not independent and identically distributed (non-IID) data, the task of training an optimal global model typically requires multiple communication rounds between the clients and the server, causing high communication overhead  [13]. Several strategies, including client selection  [28], update compression  [15], and update dropping  [39], have been proposed to reduce the communication load in each communication round. However, these strategies still require client synchronization and suffer under non-IID data distribution across clients, as local client models can drift from the global model at each communication round  [27].
$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{3}$02020404060608080100100LocalFedAvgFedCADOFedDISCOSCARClient uploaded parameters in millions (log scale)Accuracy (%)
@@
Figure 1: Uploaded parameters by each client and accuracy for various algorithms on the OpenImage dataset and ResNet-18.
@@
One-shot federated learning (OSFL) [8] offers an alternative, where the global model is learned through a single communication round between the clients and the server. OSFL can also reduce the impact of heterogeneous data distribution, as the global model is not directly formed from the local models. The single communication round can further benefit in scenarios where FL suffers from client dropout or stragglers (i.e., clients that communicate slowly) [32]. Existing OSFL approaches [36, 40, 34] rely on auxiliary dataset generation or knowledge distillation to form the global model. Knowledge distillation approaches usually require an auxiliary public dataset as the knowledge transfer medium [40]. On the other hand, dataset generation methods employ pre-trained generative models (e.g., diffusion models) to generate new data for training the global model. DMs are pre-trained with vast amounts of data, and with proper guidance, these DMs can generate realistic images that resemble a desired distribution. DMs can have an immense impact on FL, as new data resembling the clients’ distribution can be generated without access to the raw dataset of the participating clients.
@@
Current studies incorporating diffusion models in OSFL [32, 21, 34, 37] utilize classifier-guided DMs. Employing classifier-guided DMs requires auxiliary classifier training at each client, introducing computational and communication burdens. Furthermore, in some cases, the diffusion model needs to be downloaded to the clients [34]. Classifier-free DMs [12] solve these challenges by integrating the conditioning directly into the model, which is also adopted by most of the prevalent image generative models [4, 25]. Foundation models (FMs) [16, 2] can be employed for the encoding generation that can be used as conditioning without training or fine-tuning. Replacing the classifier models with encodings significantly reduces the client upload size compared to classifier-guided DM-based OSFL approaches, as shown in Fig 1. The seamless integration of pre-trained FMs and DMs in OSFL simplifies the overall framework, reduces communication load, and enhances scalability and efficiency across heterogeneous client datasets.
@@ II-A One-Shot Federated Learning
Existing OSFL approaches can be divided into two categories based on their methodology. The first category utilizes knowledge distillation to learn a global model through either data distillation [40] or model distillation [18]. In distilled one-shot federated learning (DOSFL [40], the clients share distilled synthetic data with the server, which is utilized for global model training. FedKT [18] utilizes a public auxiliary dataset and student-teacher models trained by clients to learn a global student model. The second category of methods uses auxiliary data generation at the server based on intermediary information shared by the clients. DENSE  [36] trains a generator model on local classifiers, later used to generate auxiliary data for global model training. In FedCVAE [10], the server aggregates the decoder part of the conditional variational encoders (CVAE) trained at each client and generates auxiliary data for the global model. FedDiff [21] aggregates locally trained diffusion models to form a global diffusion model for data generation. FedCADO [32] utilizes classifiers trained at each client to generate data for global model training via classifier-guided pre-trained diffusion models (DMs). FedDISC [33] utilizes data features for data generation via pre-trained DMs.
@@ II-B Federated Learning with Foundation Models
The emergence of foundation models (FMs), both large language models (LLMs) [2] and vision language models (VLMs) [16], has impacted the landscape of machine learning. The application of these FMs, however, has not yet been fully explored in FL. Yu et al.,  [35], and Charles et al.,  [5] explore training FMs in FL setups. PromptFL [9] investigates prompt learning for FMs under data scarcity in FL settings. FedDAT  [6] proposes a federated fine-tuning approach for multi-modal FMs. FedPCL [29] integrates FMs into the traditional FL process to act as class-wise prototype extractors. While FMs have the potential to mitigate data heterogeneity and communication load in FL, their full potential has not been utilized in FL settings.
@@ III-A Diffusion Models
$\begin{split}\mathcal{L}_{t}(\theta)=\mathbb{E}_{z_{0}\sim q(\mathcal{E}(x_{0})),\;\epsilon\sim\mathcal{N}(0,\;\mathbf{I}),\;t}\left[\left\|\epsilon-\hat{\epsilon}_{t}\right\|^{2}\right],\\ \hat{\epsilon}_{t}=\epsilon_{\theta}(x_{t},\;t,\;y),\end{split}$ (3)
where $\hat{\epsilon}_{t}$ provides an estimate of the score function used for data generation during the reverse process, and $\mathcal{E}(x_{0})$ denotes the encoded representation of the original input $x_{0}$.
@@ IV OSCAR: One Shot Federated Learning with Classifier-Free Diffusion Models
The traditional FL setup consists of a central server and a set of clients $\mathcal{R}$, where each client $r\in\mathcal{R}$ trains a local model $\mathbf{w}_{r}$ on its local dataset $\mathcal{D}_{r}$ in each iteration and communicates it to the server. The server is responsible for forming a global model $\mathbf{w}$ from the local client models through an aggregation function like federated averaging (FedAvg). Considering the intrinsic data heterogeneity in FL, traditional FL algorithms require multiple communication rounds between the clients and the server to form a global model, leading to significant communication overhead.
To reduce the communication load in the FL setup, we propose OSCAR, a novel one-shot federate learning (OSFL) approach. OSCAR integrates foundation and pre-trained classifier-free generative models, specifically Stable Diffusion [25], and learns a global model from the clients’ category-specific representations. OSCAR facilitates global model learning within a single communication round under a non-IID data distribution among the clients. As illustrated in Fig. 2, the OSCAR pipeline can be divided into four steps: (1) generating descriptions of the client’s client-specific data, (2) encoding features from the generated descriptions, and transmitting them to the server (3) transmitting the category-specific representations to the server, and (4) generating data on the server to facilitate final model training.
@@
Unlike existing DM-assisted OSFL approaches, OSCAR eliminates classifier training and utilizes the clients’ category-specific data features as conditioning for the DM. Each client follows a two-step approach to generate category-specific encodings. First, the client uses a VLM, specifically BLIP [17], to generate textual descriptions for all their images. Then, the client uses a text encoder, specifically CLIP [23], denoted as $\text{CLIP}_{\text{Text}}$, to generate category-specific text encodings from the textual descriptions, as shown in Eq. 6. A classifier-free diffusion model can utilize the CLIP encodings directly as text conditioning to generate new data.
@@
$x_{T-1}=\frac{1}{\sqrt{\alpha_{T}}}\left(\sqrt{\alpha_{T}}x_{T}-\hat{\epsilon}_{T}(\bar{y}_{c})\right)+\sigma_{T}\mathcal{N}(0,\mathbf{I})$ (9)
Starting from random noise $x_{T}$, the model iteratively refines the image through the sequence $x_{T}\rightarrow x_{T-1}\rightarrow\dots\rightarrow x_{0}$, guided by the category-specific encoding $\bar{y}_{c}$, to produce the final output image $x_{0}$ within the client’s distribution. In our comparisons to SOTA models, replacing the classifier model with the clients’ category-specific representations has the advantage of reducing each client’s upload size by at least 99The server generates ten images for each category-specific client representation to form a global synthesized dataset $\mathcal{D}_{syn}$, with $10\times|\mathcal{R}|\times C$ new images, where $|\mathcal{R}|$ is the number of clients and $C$ is the number of categories. As the global dataset $\mathcal{D}_{syn}$ is constructed based on the category-specific representations from individual clients, it effectively captures the heterogeneous data distribution present at each client.
@@
After generating the global synthesized dataset $\mathcal{D}_{syn}$, the server trains a centralized model $\mathbf{w}$, specifically a ResNet-18 classifier, on the synthesized dataset. This approach not only reduces the dependency on client synchronization and availability at each training and communication round but also ensures that the model can generalize well to non-IID data across clients. The server communicates the global model $\mathbf{w}$ to all clients after training, to be later used for inference locally.
@@ V Experimental Setup
  • NICO++  [38]: NICO++ contains images of size $224\times 224$ from $60$ different categories, across six domains. The dataset has two settings. In Common NICO++, all the categories share the same six domains: [autumn, dim, grass, outdoor, rock, and water]. On the other hand, the Unique NICO++ contains different domains for each category.
  • DomainNet  [22]: The DomainNet dataset contains images of 345 categories over six domains. This work uses a subset of the DomainNet dataset with 90 categories.
  • OpenImage  [14]: OpenImage is a multi-task dataset with over 1.7 million images across 600 categories. This work uses a subset of 120 categories from the dataset following the pre-processing in  [32].
@@
The data has been divided among all parties in a non-IID manner. Specifically, the data is feature distribution skewed, as each client owns data about a single domain from each category in the NICO++ and DomainNet datasets. For Openimage, the classes have been divided into six similar subgroups, where each client owns a single category from each subgroup. The number of clients is fixed to six, aligning with the number of domains in all the datasets.
@@
Local learning, federated learning, and state-of-the-art DM-assisted OSFL approaches have been considered as baselines for comparison against OSCAR. Each client trains a local standalone model on its data in local training. FedAvg [20], FedProx [19], and FedDyn [1] consider traditional FL setups with minor variations in the local objective and aggregation functions. FedDISC [33] and FedCADO [32] are DM-assisted OSFL approaches and train auxiliary models for image generation at the server.
@@ VI Results and Analysis
To facilitate direct comparison with existing baseline approaches, the classifier network in the main results reported is a ResNet-18. However, the synthesized data appears to have potential that more advanced backbones can utilize. The results for NICO++ Unique and NICO++ Common datasets with different classifier networks are reported in Table II. The results indicate that the generated data can potentially improve the global model’s optimality with an improved model architecture and may even improve more as the number of images per category increases. In the ResNet family, ResNet-101 performs the best, while the base version of the vision transformer (ViT B-16) has the best overall performance.
TABLE II: Accuracy (in %) on the test set for OSCAR with different backbone networks at the server.
@@
Number of Generated Images
This section examines the impact of the generated dataset size on the global model performance in OSCAR. Traditionally, the increase in dataset size impacts performance positively. In this case, while the initial increase in synthesized dataset size boosts the model performance, the performance remains constant, or in some cases decreases, after a certain threshold. This may indicate that data synthesized by diffusion models may best be utilized as auxiliary data for training OSFL approaches rather than as a replacement. Table III shows the result for OSCAR with varying number of samples synthesized per category of each client.
@@
Communication Analysis
As OSCAR only uploads the data encodings, it uploads the least parameters from each client. In our experiments, OSCAR uploads less than 1% of the number of parameters compared to the evaluated SOTA models. OSCAR achieves this by eliminating the classifier training, and hence classifier uploading, and each client only communicates 512 parameters for each category. Table IV shows the number of parameters each client uploads. FedCADO trains a classifier model; hence, each client uploads a model with 11.69 million parameters. While FedDISC reduces the communication size by more than 60% compared to FedCADO at each client, the upload size from each client is still more than 100 times higher than OSCAR.
@@ VII Conclusion
In this work, we propose OSCAR, a novel one-shot federated learning approach that utilizes pre-trained vision language models and classifier-free diffusion models (DMs) to train a global model in a single communication round in FL settings. OSCAR eliminates the need for training a classifier model at each client by replacing the classifier-guided DM with a classifier-free DM in the image synthesis phase. OSCAR generates category-specific data representations for each client through BLIP and CLIP foundation models, which are communicated to the server. The server generates new data samples and trains a global model on the generated data. In our experiments, OSCAR reduces the communication load by reducing the client upload size more than 100X compared to state-of-the-art DM-assisted OSFL approaches while exhibiting superior performance on four benchmarking datasets.
In future work, we want to extend OSCAR to fuse the knowledge of auxiliary generated datasets with existing learned knowledge at each client.